# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ D.1 Consistency under inner product latent space models
    $\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max(\delta,\sqrt{\frac{\log n}{n}})$
@@
  • If we have one offset covariate, the proof and the result still hold. The only difference is that we exclude on coordinate from $\beta$ in the estimation.
@@
In the node-wise estimation procedure, for any fixed node $i$, we can treat $(\hat{Z}_{i},\hat{\alpha}_{i})$ as the $\hat{\beta}$ in a logistic regression problem with true covariates $Z_{j}$, $n+1\leq j\leq n+m$, and perturbed covariates $\hat{Z}_{j},n+1\leq j\leq n+m$. We apply Lemma D.3 to all $1\leq i\leq n$ to get the result. Notice that in handling the success probability, the event involving $\hat{Z}_{i}$, $n+1\leq i\leq n+m$, is indeed shared across all $1\leq i\leq n$, for which the probability control can be improved.
@@
$H(Z_{i})=\mathbb{E}_{Z_{j},A_{ij}}\left[\nabla^{2}\ell(Z_{i};Z_{j},A_{ij})\right]=\mathbb{E}\left[\sigma^{\prime}(Z_{i}^{\top}Z_{j})Z_{j}Z_{j}^{\top}\right]$ (23)
where $\sigma^{\prime}(z)=\frac{e^{z}}{(1+e^{z})^{2}}$ is the derivative of the standard sigmoid function.
To find a uniform lower bound for the smallest eigenvalue of $H(\beta)$, it is sufficient to find a $\mu>0$ such that for any unit vector $v\in\mathbb{R}^{d}$ (i.e., $\|v\|_{2}=1$), we have $v^{\top}H(Z_{i})v\geq\mu$, for any $Z_{i}$ in the domain.
By the Cauchy-Schwarz, we have
$|Z_{i}^{\top}Z_{j}|\leq\|Z_{i}\|_{2}\|Z_{j}\|_{2}\leq R^{2}$
@@
$v^{\top}H(Z_{i})v$ $==\mathbb{E}\left[\sigma^{\prime}(Z_{i}^{\top}Z_{j})(v^{\top}Z_{j})^{2}\right]\geq w_{\min}\mathbb{E}\left[(v^{\top}Z_{j})^{2}\right]$
$w_{\min}v^{\top}\left(\mathbb{E}[Z_{j}Z_{j}^{\top}]\right)v.$ (24)
@@
Therefore, as long as we have $\Sigma\succeq\mu I$ for some $\mu>0$, the requirement hold for all $Z_{i},i\in[n]$.
@@ D.2 Consistency under the RDPG
We consider i.i.d. data $\{(x_{i},y_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{0,1\}$, generated conditional on true predictors $x_{i}$ with probability $p_{i}=\mathbb{E}[y_{i}|x_{i}]=x_{i}^{\top}\beta^{*}$, where $\beta^{*}\in\mathbb{R}^{d}$ is the true parameter vector. The squared error loss for a single data point is $\ell_{sq}(\beta;x,y)=(y-x^{\top}\beta)^{2}$.
For the proposed method, we observe perturbed predictors $\{\tilde{x}_{i}\}_{i=1}^{n}$ satisfying $\|{\tilde{x}_{i}-x_{i}}\|_{2}\leq\delta_{n}$. The average empirical losses based on true and perturbed data can hence be represented as:
$L(\beta)=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-x_{i}^{\top}\beta)^{2}.$
$\widetilde{L}(\beta)=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\tilde{x}_{i}^{\top}\beta)^{2}.$
@@
$\hat{\beta}^{\,(true)}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}L(\beta).$
$\hat{\beta}=\operatorname{argmin}\displaylimits_{\beta\in\mathbb{R}^{d}}\widetilde{L}(\beta).$
Our goal is to bound the error $\|\hat{\beta}-\beta^{*}\|_{2}$. We impose the following regularity conditions.
@@
    (Boundedness): $\|x_{i}\|_{2}\leq R_{x}$ a.s. and $\|\beta^{*}\|_{2}\leq R_{\beta}$ for constants $R_{x},R_{\beta}>0$. We also assume $0\leq x_{i}^{\top}\beta^{*}\leq 1$ for all $i$.
@@
    (Expected Design Matrix): $\Sigma_{x}=\mathbb{E}[x_{i}x_{i}^{\top}]\succeq\mu I$ for some fixed $\mu>0$. Let $\hat{\Sigma}_{x}=\frac{1}{n}\sum x_{i}x_{i}^{\top}$.
@@
    (Perturbation): $\|\tilde{x}_{i}-x_{i}\|_{2}\leq\delta_{n}$, with $\delta_{n}\to 0$ as $n\to\infty$.
@@
$\operatorname{P}\left(\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max\left\{\sqrt{\frac{d}{n}},\delta_{n}\right\}\right)\geq 1-c_{1}e^{-c_{2}n}.$
@@
Let $\Delta_{\text{total}}=\hat{\beta}-\beta^{*}$. We use the triangle inequality:
$\|\hat{\beta}-\beta^{*}\|_{2}\leq\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}+\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}.$
@@
Let $\epsilon_{i}=y_{i}-\mathbb{E}[y_{i}|x_{i}]=y_{i}-x_{i}^{\top}\beta^{*}$. $\mathbb{E}[\epsilon_{i}|x_{i}]=0$. The error of the ideal OLS estimator is given by:
$\hat{\beta}^{\,(true)}-\beta^{*}=\hat{\Sigma}_{x}^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}\right).$
Taking norms:
$\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}\leq\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}\|{\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}}\|_{2}.$
We bound the two terms on the right. Let $\epsilon_{i}=y_{i}-\mathbb{E}[y_{i}|x_{i}]=y_{i}-x_{i}^{\top}\beta^{*}$. $\mathbb{E}[\epsilon_{i}|x_{i}]=0$. The error of the ideal OLS estimator is given by:
$\hat{\beta}^{\,(true)}-\beta^{*}=\hat{\Sigma}_{x}^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}\right),$
@@
$\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}\leq\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}\|{\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}}\|_{2}.$
@@
First, the term $\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}$ involves the sample covariance concentration for i.i.d random vectors. By standard concentration result (e.g., Vershynin [2018]), we have
$\operatorname{P}\left(\left\|\hat{\Sigma}_{x}-\Sigma_{x}\right\|_{\mathrm{op}}\geq\mu/2\right)\leq 2d\exp\left(\frac{-n(\mu/2)^{2}}{8R_{x}^{4}}\right)=2d\exp\left(\frac{-n\mu^{2}}{32R_{x}^{4}}\right).$
Define $\mathcal{E}_{A}$ to be the event that $\left\|\hat{\Sigma}_{x}-\Sigma_{x}\right\|_{\mathrm{op}}\leq\mu/2$. Then $P(\mathcal{E}_{A})\geq 1-2de^{-c^{\prime}n}$ for $c^{\prime}=\mu^{2}/(32R_{x}^{4})$. On event $\mathcal{E}_{A}$, using Weyl’s inequality:
$\lambda_{\min}(\hat{\Sigma}_{x})\geq\lambda_{\min}(\Sigma_{x})-\left\|\hat{\Sigma}_{x}-\Sigma_{x}\right\|_{\mathrm{op}}\geq\mu-\mu/2=\mu/2,$
@@
$\left\|\hat{\Sigma}_{x}^{-1}\right\|_{\mathrm{op}}\leq\frac{1}{\lambda_{\min}(\hat{\Sigma}_{x})}\leq\frac{2}{\mu}.$
For the second term, let $S=\frac{1}{n}\sum Z_{i}$ where $Z_{i}=x_{i}\epsilon_{i}$. As noted, $\mathbb{E}[Z_{i}]=0$ and $|\epsilon_{i}|\leq 1$. Thus, $\|Z_{i}\|_{2}=\|x_{i}\epsilon_{i}\|_{2}\leq\|x_{i}\|_{2}|\epsilon_{i}|\leq R_{x}$. The vectors $Z_{i}$ are bounded, independent, mean-zero random vectors. By Hoeffding inequality on bounded random vectors, we have
$\mathbb{P}(\|{\frac{1}{n}\sum_{i=1}^{n}x_{i}\epsilon_{i}}\|_{2}\leq C_{S}\sqrt{\frac{d}{n}})\geq 1-c_{1}^{\prime}e^{-c_{2}^{\prime}n}.$
@@
$P(\mathcal{E}_{B})\geq 1-c^{\prime}_{1}e^{-c^{\prime}_{2}n}.$ (25)
Let $\mathcal{E}_{true}=\mathcal{E}_{A}\cap\mathcal{E}_{B}$. Then $P(\mathcal{E}_{true})\geq 1-c^{\prime\prime}_{1}e^{-c^{\prime\prime}_{2}n}$. On event $\mathcal{E}_{true}$: $\|\hat{\beta}^{\,(true)}-\beta^{*}\|_{2}\leq C_{true}\sqrt{d/n}$ for constant $C_{true}>0$.
Next, we bound $\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}$.
Let $\Delta=\hat{\beta}-\hat{\beta}^{\,(true)}$. The first-order condition for $\hat{\beta}$ is $\nabla\widetilde{L}(\hat{\beta})=0$. The Hessian $\nabla^{2}\widetilde{L}(\beta)=2\hat{\Sigma}_{\tilde{x}}=\frac{2}{n}\sum\tilde{x}_{i}\tilde{x}_{i}^{\top}$ is constant w.r.t. $\beta$. Since $L$ is quadratic, we have
$0=\nabla\widetilde{L}(\hat{\beta})=\nabla\widetilde{L}(\hat{\beta}^{\,(true)})+2\hat{\Sigma}_{\tilde{x}}\Delta$
Substituting $\nabla L(\hat{\beta}^{\,(true)})=0$ into the equation above, we have
$\nabla\widetilde{L}(\hat{\beta}^{\,(true)})=\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)}).$
@@
$\Delta=-\frac{1}{2}\hat{\Sigma}_{\tilde{x}}^{-1}\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})$
@@
$\|\Delta\|_{2}\leq\frac{1}{2}\left\|\hat{\Sigma}_{\tilde{x}}^{-1}\right\|_{\mathrm{op}}\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}.$
@@
$\left\|\hat{\Sigma}_{\tilde{x}}-\hat{\Sigma}_{x}\right\|_{\mathrm{op}}\leq\frac{1}{n}\sum\left\|\tilde{x}_{i}\tilde{x}_{i}^{\top}-x_{i}x_{i}^{\top}\right\|_{\mathrm{op}}\leq(2R_{x}+\delta_{n})\delta_{n}.$
@@
$\lambda_{\min}(\hat{\Sigma}_{\tilde{x}})\geq\mu/4.$
Let $\mathcal{E}_{C}$ be this event ($P(\mathcal{E}_{C})\geq P(\mathcal{E}_{A})$).
For the second term of gradient difference: $\nabla(\widetilde{L}-L)(\beta)=\frac{1}{n}\sum T_{i}(\beta)$, where
@@
Thus, on event $\mathcal{E}_{true}$, we have
$\|T_{i}(\hat{\beta}^{\,(true)})\|_{2}\leq(|\tilde{x}_{i}^{\top}\hat{\beta}^{\,(true)}|+|y_{i}|)\delta_{n}+|(\tilde{x}_{i}-x_{i})^{\top}\hat{\beta}^{\,(true)}|\|x_{i}\|_{2}=O(\delta_{n}).$
@@
$\|\nabla(\widetilde{L}-L)(\hat{\beta}^{\,(true)})\|_{2}\leq\frac{1}{n}\sum\|T_{i}(\hat{\beta}^{\,(true)})\|_{2}\leq O(\delta_{n}).$
Substituting these bounds into the inequality for $\|\Delta\|_{2}$, we get on event $\mathcal{E}=\mathcal{E}_{true}\cap\mathcal{E}_{C}$:
$\|\hat{\beta}-\hat{\beta}^{\,(true)}\|_{2}=\|\Delta\|_{2}\leq\frac{1}{2}\left(\frac{4}{\mu}\right)(C_{G}\delta_{n})=\frac{2C_{G}}{\mu}\delta_{n}.$
Finally, let $\mathcal{E}=\mathcal{E}_{true}\cap\mathcal{E}_{C}$. Taking $d$ as a constant, using the two bounds, we get
$\|\hat{\beta}-\beta^{*}\|_{2}\leq C\max\left\{\sqrt{\frac{1}{n}},\delta_{n}\right\}$
@@
Under the random dot product graph model, assume that the latent random vector $Z_{i}$’s are bounded. Suppose $\mathbb{E}[Z_{i}Z_{i}^{T}]\succeq\mu I_{d}$ for some $\mu>0$. Moreover, on the hold-out data, suppose that with probability tending to 1, we can achieve

### v3

@@ D.1 Consistency under inner product latent space models
    $\|\hat{\beta}-\beta^{*}\|\leq C\max\left\{\delta,\sqrt{\frac{\log n}{n}}\right\}$
@@
  • If we have one offset covariate, the proof and the result still hold. The only difference is that we exclude one coordinate from $\beta$ in the estimation.
@@
In the node-wise estimation procedure, for any fixed node $i$, we can treat $(\hat{X}_{i},\hat{\alpha}_{i})$ as the $\hat{\beta}$ in a logistic regression problem with true covariates $Z_{j}$, $n+1\leq j\leq n+m$, and perturbed covariates $\hat{Z}_{j}$, $n+1\leq j\leq n+m$. We apply Lemma D.3 to all $1\leq i\leq n$ to get the result. Notice that in handling the success probability, the event involving $\hat{Z}_{i}$, $n+1\leq i\leq n+m$, is indeed shared across all $1\leq i\leq n$, for which the probability control can be improved.
@@
$H(Z_{i})=\mathbb{E}_{Z_{j},A_{ij}}[\nabla^{2}\ell(Z_{i};Z_{j},A_{ij})]=\mathbb{E}[\sigma^{\prime}(Z_{i}^{\top}Z_{j})Z_{j}Z_{j}^{\top}]$ (23)
where $\sigma^{\prime}(z)=e^{z}/(1+e^{z})^{2}$ is the derivative of the standard sigmoid function.
To find a uniform lower bound for the smallest eigenvalue of $H(\beta)$, it is sufficient to find a $\mu>0$ such that for any unit vector $v\in\mathbb{R}^{d}$ (i.e., $\|v\|=1$), we have $v^{\top}H(Z_{i})v\geq\mu$, for any $Z_{i}$ in the domain.
By the Cauchy-Schwarz inequality, we have
$|Z_{i}^{\top}Z_{j}|\leq\|Z_{i}\|\|Z_{j}\|\leq R^{2}$
@@
$v^{\top}H(Z_{i})v$ $=\mathbb{E}[\sigma^{\prime}(Z_{i}^{\top}Z_{j})(v^{\top}Z_{j})^{2}]\geq w_{\min}\mathbb{E}[(v^{\top}Z_{j})^{2}]=w_{\min}v^{\top}(\mathbb{E}[Z_{j}Z_{j}^{\top}])v.$ (24)
@@
Therefore, as long as we have $\Sigma\succeq\mu I_{d}$ for some $\mu>0$, the requirement hold for all $Z_{i}$, $i\in[n]$.
@@ D.2 Consistency under the RDPG
We consider i.i.d. data $\{(x_{i},y_{i})\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{0,1\}$, generated conditional on true predictors $x_{i}$ with probability $p_{i}=\mathbb{E}[y_{i}\,|\,x_{i}]=x_{i}^{\top}\beta^{*}$, where $\beta^{*}\in\mathbb{R}^{d}$ is the true parameter vector. The squared error loss for a single data point is $\ell_{\rm sq}(\beta;x,y)=(y-x^{\top}\beta)^{2}$.
For the proposed method, we observe perturbed predictors $\{\tilde{x}_{i}\}_{i=1}^{n}$ satisfying $\|{\tilde{x}_{i}-x_{i}}\|\leq\delta_{n}$. The average empirical losses based on true and perturbed data can hence be represented as:
$\mathcal{L}(\beta)=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-x_{i}^{\top}\beta)^{2}.$
$\tilde{\mathcal{L}}(\beta)=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\tilde{x}_{i}^{\top}\beta)^{2}.$
@@
$\hat{\beta}_{\rm true}=\mathop{\arg\min}_{\beta\in\mathbb{R}^{d}}\mathcal{L}(\beta).$
$\hat{\beta}=\mathop{\arg\min}_{\beta\in\mathbb{R}^{d}}\tilde{\mathcal{L}}(\beta).$
Our goal is to bound the error $\|\hat{\beta}-\beta^{*}\|$. We impose the following regularity conditions.
@@
    (Boundedness): $\|x_{i}\|\leq R_{x}$ a.s. and $\|\beta^{*}\|\leq R_{\beta}$ for constants $R_{x},R_{\beta}>0$. We also assume $0\leq x_{i}^{\top}\beta^{*}\leq 1$ for all $i$.
@@
    (Expected Design Matrix): $\Sigma_{x}=\mathbb{E}[x_{i}x_{i}^{\top}]\succeq\mu I_{d}$ for some fixed $\mu>0$. Let $\hat{\Sigma}_{x}=\frac{1}{n}\sum x_{i}x_{i}^{\top}$.
@@
    (Perturbation): $\|\tilde{x}_{i}-x_{i}\|\leq\delta_{n}$, with $\delta_{n}\to 0$ as $n\to\infty$.
@@
$\mathbb{P}\left(\|\hat{\beta}-\beta^{*}\|\leq C\max\left\{\sqrt{\frac{d}{n}},\delta_{n}\right\}\right)\geq 1-c_{1}e^{-c_{2}n}.$
@@
Let $\Delta_{\mathrm{total}}=\hat{\beta}-\beta^{*}$. We use the triangle inequality:
$\|\hat{\beta}-\beta^{*}\|\leq\|\hat{\beta}-\hat{\beta}_{\rm true}\|+\|\hat{\beta}_{\rm true}-\beta^{*}\|.$
@@
Let $\varepsilon_{i}=y_{i}-\mathbb{E}[y_{i}\,|\,x_{i}]=y_{i}-x_{i}^{\top}\beta^{*}$. $\mathbb{E}[\varepsilon_{i}\,|\,x_{i}]=0$. The error of the ideal OLS estimator is given by:
$\hat{\beta}_{\rm true}-\beta^{*}=\hat{\Sigma}_{x}^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}\varepsilon_{i}\right),$
@@
$\|\hat{\beta}_{\rm true}-\beta^{*}\|\leq\|\hat{\Sigma}_{x}^{-1}\|_{\rm op}\left\|\frac{1}{n}\sum_{i=1}^{n}x_{i}\varepsilon_{i}\right\|.$
@@
First, the term $\|\hat{\Sigma}_{x}^{-1}\|_{\rm op}$ involves the sample covariance concentration for i.i.d. random vectors. By standard concentration result (e.g., Vershynin [2018]), we have
$\mathbb{P}\left(\|\hat{\Sigma}_{x}-\Sigma_{x}\|_{\rm op}\geq\frac{\mu}{2}\right)\leq 2d\exp\left(\frac{-n(\mu/2)^{2}}{8R_{x}^{4}}\right)=2d\exp\left(\frac{-n\mu^{2}}{32R_{x}^{4}}\right).$
Define $\mathcal{E}_{A}$ to be the event that $\|\hat{\Sigma}_{x}-\Sigma_{x}\|_{\rm op}\leq\mu/2$. Then $\mathbb{P}(\mathcal{E}_{A})\geq 1-2de^{-c^{\prime}n}$ for $c^{\prime}=\mu^{2}/(32R_{x}^{4})$. On event $\mathcal{E}_{A}$, using Weyl’s inequality:
$\lambda_{\min}(\hat{\Sigma}_{x})\geq\lambda_{\min}(\Sigma_{x})-\left\|\hat{\Sigma}_{x}-\Sigma_{x}\right\|_{\mathrm{op}}\geq\mu-\frac{\mu}{2}=\frac{\mu}{2},$
@@
$\|\hat{\Sigma}_{x}^{-1}\|_{\rm op}\leq\frac{1}{\lambda_{\min}(\hat{\Sigma}_{x})}\leq\frac{2}{\mu}.$
For the second term, let $S=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ where $Z_{i}=x_{i}\varepsilon_{i}$. As noted, $\mathbb{E}[Z_{i}]=0$ and $|\varepsilon_{i}|\leq 1$. Thus, $\|Z_{i}\|=\|x_{i}\varepsilon_{i}\|\leq\|x_{i}\||\varepsilon_{i}|\leq R_{x}$. The vectors $Z_{i}$ are bounded, independent, mean-zero random vectors. By Hoeffding’s inequality on bounded random vectors, we have
$\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}x_{i}\varepsilon_{i}\right\|\leq C_{S}\sqrt{\frac{d}{n}}\right)\geq 1-c_{1}^{\prime}e^{-c_{2}^{\prime}n}.$
@@
$\mathbb{P}(\mathcal{E}_{B})\geq 1-c^{\prime}_{1}e^{-c^{\prime}_{2}n}.$ (25)
Let $\mathcal{E}_{\rm true}=\mathcal{E}_{A}\cap\mathcal{E}_{B}$. Then $\mathbb{P}(\mathcal{E}_{\rm true})\geq 1-c^{\prime\prime}_{1}e^{-c^{\prime\prime}_{2}n}$. On event $\mathcal{E}_{\rm true}$: $\|\hat{\beta}_{\rm true}-\beta^{*}\|\leq C_{\rm true}\sqrt{d/n}$ for constant $C_{\rm true}>0$.
Next, we bound $\|\hat{\beta}-\hat{\beta}_{\rm true}\|$.
Let $\Delta=\hat{\beta}-\hat{\beta}_{\rm true}$. The first-order condition for $\hat{\beta}$ is $\nabla\tilde{\mathcal{L}}(\hat{\beta})=0$. The Hessian $\nabla^{2}\tilde{\mathcal{L}}(\beta)=2\hat{\Sigma}_{\tilde{x}}=\frac{2}{n}\sum\tilde{x}_{i}\tilde{x}_{i}^{\top}$ is constant w.r.t. $\beta$. Since $L$ is quadratic, we have
$0=\nabla\tilde{\mathcal{L}}(\hat{\beta})=\nabla\tilde{\mathcal{L}}(\hat{\beta}_{\rm true})+2\hat{\Sigma}_{\tilde{x}}\Delta$
Substituting $\nabla\mathcal{L}(\hat{\beta}_{\rm true})=0$ into the equation above, we have
$\nabla\tilde{\mathcal{L}}(\hat{\beta}_{\rm true})=\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true}).$
@@
$\Delta=-\frac{1}{2}\hat{\Sigma}_{\tilde{x}}^{-1}\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})$
@@
$\|\Delta\|\leq\frac{1}{2}\|\hat{\Sigma}_{\tilde{x}}^{-1}\|_{\rm op}\|\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})\|.$
@@
$\|\hat{\Sigma}_{\tilde{x}}-\hat{\Sigma}_{x}\|_{\rm op}\leq\frac{1}{n}\sum_{i=1}^{n}\|\tilde{x}_{i}\tilde{x}_{i}^{\top}-x_{i}x_{i}^{\top}\|_{\rm op}\leq(2R_{x}+\delta_{n})\delta_{n}.$
@@
$\lambda_{\min}(\hat{\Sigma}_{\tilde{x}})\geq\frac{\mu}{4}.$
Let $\mathcal{E}_{C}$ be this event ($\mathbb{P}(\mathcal{E}_{C})\geq\mathbb{P}(\mathcal{E}_{A})$).
For the second term of gradient difference: $\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\beta)=\frac{1}{n}\sum_{i=1}^{n}T_{i}(\beta)$, where
@@
Thus, on event $\mathcal{E}_{\rm true}$, we have
$\|T_{i}(\hat{\beta}_{\rm true})\|\leq(|\tilde{x}_{i}^{\top}\hat{\beta}_{\rm true}|+|y_{i}|)\delta_{n}+|(\tilde{x}_{i}-x_{i})^{\top}\hat{\beta}_{\rm true}|\|x_{i}\|=\mathcal{O}(\delta_{n}).$
@@
$\|\nabla(\tilde{\mathcal{L}}-\mathcal{L})(\hat{\beta}_{\rm true})\|\leq\frac{1}{n}\sum_{i=1}^{n}\|T_{i}(\hat{\beta}_{\rm true})\|\leq\mathcal{O}(\delta_{n}).$
Substituting these bounds into the inequality for $\|\Delta\|$, we get on event $\mathcal{E}=\mathcal{E}_{\rm true}\cap\mathcal{E}_{C}$:
$\|\hat{\beta}-\hat{\beta}_{\rm true}\|=\|\Delta\|\leq\frac{1}{2}\left(\frac{4}{\mu}\right)(C_{G}\delta_{n})=\frac{2C_{G}}{\mu}\delta_{n}.$
Finally, let $\mathcal{E}=\mathcal{E}_{\rm true}\cap\mathcal{E}_{C}$. Taking $d$ as a constant, using the two bounds, we get
$\|\hat{\beta}-\beta^{*}\|\leq C\max\left\{\sqrt{\frac{1}{n}},\delta_{n}\right\}$
@@
Under the random dot product graph model, assume that the latent random vector $Z_{i}$’s are bounded. Suppose $\mathbb{E}[Z_{i}Z_{i}^{\top}]\succeq\mu I_{d}$ for some $\mu>0$. Moreover, on the hold-out data, suppose that with probability tending to 1, we can achieve